Use the Principle of Mathematical Induction to prove that "((2n) !)\/(2^n n !)" is odd for all positive integers.


1
Expert's answer
2022-01-05T04:57:28-0500

Let us use the Principle of Mathematical Induction to prove that "\frac{(2n) !}{2^n n !}" is odd for all positive integers.

For "n=1" we get that "\frac{2 !}{2 \cdot 1 !}=1" is odd.

Assume that the statement is true for "n=k," that is "\frac{(2k)!}{2^k k!}" is odd.

Let us prove for "n=k+1."

It follows that "\frac{(2(k+1))!}{2^{k+1} (k+1)!}=\frac{(2k)!(2k+2)(2k+1)}{2^k2(k+1) k!}
=\frac{(2k)!}{2^k k!}(2k+1)." Taking into account that "\frac{(2k)!}{2^k k!}" and "2k+1" are odd, we conclude that "\frac{(2k)!}{2^k k!}(2k+1)" is also odd. Therefore, we proved that "\frac{(2(k+1))!}{2^{k+1} (k+1)!}" is odd.

We conclude that by the Principle of Mathematical Induction  "\frac{(2n) !}{2^n n !}" is odd for all positive integers.